Let A be any n x n matrix and let P be an n x n orthogonal matrix. Prove that jAj D jP


APj:


1
Expert's answer
2021-05-19T17:25:13-0400

A is "n\times n" symmetric matrix.

its all eigen values are greater than 0.


Characteristics polynomial is "det(A-\lambda I_n)=0"


"\lambda" are the eigenvalues of A.

As P is orthogonal matrix so "P^T=P^{-1}"


Characteristics polynomial for "P^TAP" is given as-


"det(P^TAP-\lambda I_n)


\\
=det(P^{-1}AP-\lambda P^{-1}P)


\\
=det{P^{-1}(A-\lambda I_n)P}


\\
=det(P^{-1})det (P)det(A-\lambda I_n)


\\
=det(P^{-1}P)det(A-\lambda I_n)
\\


=det(A-\lambda I_n)"



"(P^TAP)^T=(PA)^T(P^T)^T=P^TA^TP=P^TAP, as A^T=A"


So "P^TAP" is symmetric matrix.


From characteristics equation, eigenvalues of "P^TAP" is the same as eigenvalues of A.